OpenStudy (anonymous):

Solve cot^2 x -csc x -1=0

OpenStudy (anonymous):

sin^2 +cos^2 =1 (sin^2)/(sin^2) + (cos^2)/(sin^2) =1/(sin^2) cot^2=csc^2-1 cot^2 x -csc x -1=0 csc^2-1 -csc x -1=0 csc^2 -csc -2=0 (cscx-2)(cscx+1)=0 cscx =2 (rej) or cscx=-1 x can then be solved